OpenStudy (anonymous):

Find the equation of the circle touching the x-axis and passing through the points (3,1) and (10,8). Thanks.

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2

OpenStudy (dumbcow):

yep center; (h,k)

OpenStudy (anonymous):

then what's next!?

OpenStudy (anonymous):

then what's next!?

myininaya (myininaya):

PLUG in (3,1) (3-h)^2+(1-k)^2=r^2 plug in (10,8) (10-h)^2+(8-k)^2=r^2 touching x-axis means we also have (x_0,0) (x_0-h)^2+(0-k)^2=r^2

OpenStudy (anonymous):

then what's next!?

OpenStudy (anonymous):

last one of these was not so easy. i bet this one is not also

OpenStudy (dumbcow):

find the center: it will be equal distance from each point use distance formula

OpenStudy (anonymous):

try to figure out the radius given the points

OpenStudy (anonymous):

i can't wait to see this. easy to say, not so easy to do

OpenStudy (anonymous):

yea it is tricky! :p

OpenStudy (anonymous):

You sure one of those isn't a center point? :(

myininaya (myininaya):

so far i have gotten \[h+k=11, k=\frac{9 \pm \sqrt{2r^2-49}}{2}, h=11-\frac{9 \pm \sqrt{2r^2-49}}{2}\]

OpenStudy (anonymous):

i have an idea! (probably wrong) put the center at (h,k) and point of tangency is (0,k) . distance from (9,k) to (h,k) is h

OpenStudy (anonymous):

I can't see geometrically how it is possible to have both points in the circle and have the circle touch the - or + parts of the x-axis. In that case (3,1) would have to be part of a line segment that represents a chord in the circle. Now, you can make a circle that countains the two points and all you have to do is find the distance between the two points and then use the midpoint as your center (h,k) in the equation the tother member mentioned.

OpenStudy (dumbcow):

use the equations myininaya did above. set them equal to each other, since r^2 = r^2 Then solve for h: \[(3-h)^{2}+(1-k)^{2} = (10-h)^{2} +(8-k)^{2}\] \[\rightarrow h = -k+11\] Because its resting on x-axis, x_0 = h and r=k \[h = 11-r\] \[\rightarrow (3-(11-r))^{2} + (1-r)^{2} = r^{2}\] solve for r \[(r^{2}-16r+64)+(r^{2}-2r+1) =r^{2}\] \[r^{2} -18r +65 = 0\] \[(r-13)(r-5) = 0\] \[r=5, h=6, k =5\]

OpenStudy (anonymous):

remember teh distance is given by the formula

OpenStudy (anonymous):

we have \[h^2=(h-10)^2+(k-8)^2 = (h-3)^2+(k-1)^2\] \[h^2=h^2-20h+100+k^2-16k+64\] \[-20h +k^2-16k+164=0\]

OpenStudy (anonymous):

oh i am too late. probably wouldn't have worked in any case!

OpenStudy (dumbcow):

Equation: \[(x-6)^{2} + (y-5)^{2} = 25\]

OpenStudy (anonymous):

especially since i had it tangent to the y - axis (doh)

OpenStudy (dumbcow):

satellite, yeah i think you had the right idea, just switched the axis

OpenStudy (anonymous):

yeah no wonder i had a hard time drawing it!

OpenStudy (dumbcow):

no because its resting on x-axis Not y-axis

myininaya (myininaya):

oh oops

OpenStudy (anonymous):

i am going to try to redo this the way i did it before, see if i come up with an answer.

OpenStudy (anonymous):

there were so many different circles that could have crossed those points!!

OpenStudy (anonymous):

r = k is the kicker.

OpenStudy (dumbcow):

yeah i don't think you can do it unless you know its touching an axis

myininaya (myininaya):

fantastic cow

OpenStudy (anonymous):

yeah this is really nice